OpenStudy (anonymous):

HELPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPP :) The lengths of the sides of a triangle are 3x, 5x -12 and x + 20. Find the value of x so that the triangle is isoceles.

OpenStudy (anonymous):

just set two of the equations equal to each other and solve for x for example\[3x=x+10\]

OpenStudy (anonymous):

So would i do 3x = x + 20?

OpenStudy (anonymous):

get all your xs on one side \[3x=x+20\]\[-x = -x\]\[2x=20\] and what do you do next?

OpenStudy (anonymous):

you divide 2x on each side and you get x = 10 right?

OpenStudy (anonymous):

yes you got it

OpenStudy (anonymous):

okay that would be the answer?

OpenStudy (anonymous):

yes by the way you wrote the question x=10 would be the answer

OpenStudy (anonymous):

Lol that was simple thank you!

OpenStudy (anonymous):

np